# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ Abstract
Mixture of Experts (MoE) models are highly effective in scaling model capacity while preserving computational efficiency, with the gating network, or router, playing a central role by directing inputs to the appropriate experts. In this paper, we establish a novel connection between MoE frameworks and attention mechanisms, demonstrating how quadratic gating can serve as a more expressive and efficient alternative. Motivated by this insight, we explore the implementation of quadratic gating within MoE models, identifying a connection between the self-attention mechanism and the quadratic gating. We conduct a comprehensive theoretical analysis of the quadratic softmax gating MoE framework, showing improved sample efficiency in expert and parameter estimation. Our analysis provides key insights into optimal designs for quadratic gating and expert functions, further elucidating the principles behind widely used attention mechanisms. Through extensive evaluations, we demonstrate that the quadratic gating MoE outperforms the traditional linear gating MoE. Moreover, our theoretical insights have guided the development of a novel attention mechanism, which we validated through extensive experiments. The results demonstrate its favorable performance over conventional models across various tasks.
@@ 1 Introduction
The mixture of experts (MoE) [15, 17] architecture has recently become a powerful approach in conditional computation, driving many advances in machine learning. Unlike dense models, MoEs dynamically activates only a subset of network, known as “expert”, for each input. This results in an efficient form of conditional computation. A notable modern example is the sparsely gated MoE [38], which significantly increases the model capacity without a corresponding increase in computational costs [19, 10, 11, 31]. This has made MoEs crucial for scaling up large language [16, 32, 44, 8], vision [34, 20, 35], and multimodal models [24, 12], resulting in outstanding performance in different areas.
The mixture of experts (MoE) model includes $N$ expert networks and a gating network. Each expert ${\bm{h}}_{i}(\cdot;{\bm{\eta}}_{i})$ transforms the input ${\bm{x}}\in\mathbb{R}^{d}$ into an output in $\mathbb{R}^{d_{\rm o}}$. The gating network computes the gating probabilities ${\bm{g}}({\bm{x}};\bm{\Theta}_{g})=(g_{1}({\bm{x}};\bm{\Theta}_{g}),\ldots,g_{N}({\bm{x}};\bm{\Theta}_{g}))$, with $\bm{\Theta}_{g}$ as learnable parameters. The model output ${\bm{y}}$ is the weighted sum of expert outputs: ${\bm{y}}=\sum_{i=1}^{N}g_{i}({\bm{x}};\bm{\Theta}_{g})\cdot{\bm{h}}_{i}({\bm{x}};{\bm{\eta}}_{i})$.
In the original MoE framework [15, 17], the gating network ${\bm{g}}$ computes scores for each expert using a parameterized network ${\bm{s}}({\bm{x}};\bm{\Theta}_{g})=\left(s_{1}({\bm{x}};{\bm{\theta}}_{1}),\ldots,s_{N}({\bm{x}};{\bm{\theta}}_{N})\right)\in\mathbb{R}^{N}$ and uses a softmax function to normalize these scores into a gating probability vector. Consequently, the scalar $g_{i}({\bm{x}};\bm{\Theta}_{g})$ can be expressed as
$g_{i}({\bm{x}};\bm{\Theta}_{g})=\frac{\exp{\left(s_{i}({\bm{x}};\bm{\Theta}_{g})\right)}}{\sum_{j=1}^{N}\exp{\left(s_{i}({\bm{x}};\bm{\Theta}_{g})\right)}},\quad i=1,\ldots,N.$ (1)
The scoring network typically utilizes a linear mapping, which has been consistently adopted in modern sparse MoEs [38, 19, 10, 16, 6, 25] due to its simplicity and scalability.
In this paper, we initially establish a connection between the MoE framework and the attention mechanism. In particular, we discuss that both MoE models and attention mechanisms aim to allocate computational resources efficiently by prioritizing relevant aspects of the input data. MoE models achieve this by using a gating network to selectively activate specialized experts, while the attention mechanism computes attention weights to emphasize important tokens or features in context. A detailed discussion is provided in Section 2.3. Additionally, we argue that the self-attention mechanism [41] assesses the interactions between different parts of the input to assign attention weights, which can be linked to the use of a quadratic scoring function in the gating network.
Inspired by this connection, we consider a class of quadratic gating MoE [21] models as an alternative to the conventional linear gating model. By allowing for more flexible decision boundaries through a quadratic scoring function, these models aim to provide better adaptability in expert selection. We particularly focus on two quadratic gatings: (1) Quadratic polynomial gating, which is inspired by the linear embeddings with biases of keys and queries in the attention (see Section 2.3) and given by:
$g_{i}({\bm{x}};\bm{\Theta}_{g})=\frac{\exp\left({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+{\bm{b}}_{i}^{\top}{\bm{x}}+c_{i}\right)}{\sum_{j=1}^{N}\exp\left({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+{\bm{b}}_{j}^{\top}{\bm{x}}+c_{j}\right)},$ (2)
where $\bm{\Theta}_{g}=\{({\bm{A}}_{i},{\bm{b}}_{i},c_{i})\in\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R},i=1,\ldots,N\}$ is the set of learnable parameters; (2) Quadratic monomial gating, which is motivated by the widely used linear embeddings without biases of keys and queries in the attention (see Section 2.3) and given by:
$g_{i}({\bm{x}};\bm{\Theta}_{g})=\frac{\exp\left({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+c_{i}\right)}{\sum_{j=1}^{N}\exp\left({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+c_{j}\right)},$ (3)
where $\bm{\Theta}_{g}=\{({\bm{A}}_{i},c_{i})\in\mathbb{R}^{d\times d}\times\mathbb{R},i=1,\ldots,N\}$ is the set of learnable parameters. In Appendix A, we further provided related literature on these choices of quadratic gatings.
@@
Table 1: Summary of parameter estimation rates under the mixture of strongly identifiable experts models equipped with the quadratic polynomial gate and the quadratic monomial gate.
Gate Loss $\exp(c^{*}_{j})$ ${\bm{A}}^{*}_{j}$ ${\bm{b}}^{*}_{j}$ ${\bm{\eta}}^{*}_{j}$
Quadratic Polynomial (Thm. 3.3) $\mathcal{L}_{1}$ $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{A}_{j}|)})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/2\bar{r}(|\mathcal{A}_{j}|)})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$
Quadratic Monomial (Thm. B.3) $\mathcal{L}_{3}$ $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$
Contributions. The paper has three main contributions, which can be summarized as follows:
1. Connection between attention mechanism and quadratic gating MoE. We begin by rigorously defining the MoE framework and the attention mechanism, highlighting the similarity in their formulations (see Definitions 2.1 and  2.2). Building on the similarity in their formulations, we unify these two popular concepts by introducing the Attention Gating Mixture of Experts framework (see Definition 2.3). This framework specifically highlights that the MoE framework can be effectively formulated as an attention mechanism. In particular, we connect the concepts of query, keys, and values from the attention mechanism to the MoE framework, creating a cohesive and robust unified model. We further demonstrate that the quadratic gating MoE is intrinsically linked to the self-attention mechanism, illustrating that it can be viewed as a special case within this unified framework (see Equation 8). This connection motivates us to study the quadratic gating MoE in more detail.
@@
2. Theoretical analysis of the quadratic gating MoE. We explore the effects of two variants of the quadratic gating, namely the quadratic polynomial gating and the quadratic monomial gating, on the convergence of parameter and expert estimation under the MoE models. For the convergence analysis of each gate, we provide a corresponding strong identifiability condition (see Definitions 3.2 and B.2) to characterize the compatible structure of experts with that gate. Based on those conditions, we show that experts formulated as neural networks with activation functions such as $\mathrm{ReLU}$ and $\tanh$ require fewer data to approximate than linear experts (see Theorems 3.3 and C.1).
@@
3. Practical implications. The convergence analysis of the quadratic gating MoE models provides two practical implications. Firstly, they show the benefits of quadratic monomial gating over quadratic polynomial gating, which confirms the benefits of the widely used linear embeddings of the keys and queries without bias terms in the attention mechanism in practice. Secondly, the theoretical analysis encourages the usage of non-linear experts over linear experts in quadratic gating MoE models. Given that insight, we propose a novel active-attention mechanism in equation (20) by replacing the linear value matrix by the non-linear value matrix in the attention mechanism. Through extensive empirical evaluation, we show the favorable performance of the proposed active-attention over standard attention in various tasks.
@@
Organization. The paper proceeds as follows. In Section 2, we provide background on MoE layer, attention mechanism and their connection. Next, we investigate the convergence behavior of parameter and expert estimation under the MoE model with two variants of the quadratic gate, and present two important practical implications from that analysis in Section 3. Then, we highlight some practical implications from our theory in Section 4. Subsequently, we conduct extensive experiments to empirically justify the theoretical results and favorable performance of the active-attention mechanism in Section 5. Finally, we conclude the paper in Section 6. Full proofs and the remaining materials are deferred to the Appendices.

### v3

# Quadratic Gating Mixture of Experts: Statistical Insights into Self-Attention

@@ Abstract
Mixture of Experts (MoE) models are well known for effectively scaling model capacity while preserving computational overheads. In this paper, we establish a rigorous relation between MoE and the self-attention mechanism, showing that each row of a self-attention matrix can be written as a quadratic gating mixture of linear experts. Motivated by this connection, we conduct a comprehensive convergence analysis of MoE models with two different quadratic gating functions, namely the quadratic polynomial gate and the quadratic monomial gate, offering useful insights into the design of gating and experts for the MoE framework. First, our analysis indicates that the use of the quadratic monomial gate yields an improved sample efficiency for estimating parameters and experts compared to the quadratic polynomial gate. Second, parameter and expert estimation rates become significantly faster when employing non-linear experts in place of linear experts. Combining these theoretical insights with the above link between MoE and self-attention, we propose a novel active-attention mechanism where we apply a non-linear activation function to the value matrix in the formula of self-attention. Finally, we demonstrate that the proposed active-attention outperforms the standard self-attention through several extensive experiments in various tasks, including image classification, language modeling, and multivariate time series forecasting.
@@ 1 Introduction
Mixture of experts (MoE) is a statistical machine learning that aggregates the power of multiple specialized sub-models, known as experts, based on the principle of divide and conquer. These experts can be formulated as classifiers chen2022theory ; nguyen2024general , regression models faria2010regression ; kwon_em_2020 , or feed-forward networks (FFNs) shazeer2017topk ; dai2024deepseekmoe , which we will denote by $\mathcal{E}_{j}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$, for $1\leq j\leq N$, where $N\in\mathbb{N}$ stands for the number of experts. Another key component of MoE is a gating network, which is in charge of allocating input-dependent softmax weight to each expert. In particular, given an input ${\bm{x}}\in\mathbb{R}^{d}$, the gating network first computes expert affinity scores that quantify the similarity between the expert specialization and the input domain, denoted by ${\bm{s}}({\bm{x}}):=(s_{1}({\bm{x}}),s_{2}({\bm{x}}),\ldots,s_{N}({\bm{x}}))\in\mathbb{R}^{N}$. Next, the gating values (mixture weights) are calculated as the softmax normalization of these affinity scores, that is, $\mathcal{G}({\bm{x}})_{j}=\mathrm{Softmax}({\bm{s}}({\bm{x}}))_{j}:=\frac{\exp(s_{j}({\bm{x}}))}{\sum_{i=1}^{N}\exp(s_{i}({\bm{x}}))}$, for $1\leq j\leq N$. As a result, the gating network can be trained to assign higher weights to more relevant experts. This property helps MoE become more adaptive and obtain a greater expressive power than conventional mixture models Lindsay-1995 . Then, the final output of MoE is represented as an adaptively weighted sum of the expert outputs:
${\bm{y}}=\sum_{j=1}^{N}\mathcal{G}({\bm{x}})_{j}\cdot\mathcal{E}_{j}({\bm{x}})=\sum_{j=1}^{N}\mathrm{Softmax}({\bm{s}}({\bm{x}}))_{j}\cdot\mathcal{E}_{j}({\bm{x}}).$ (1)
Thanks to the adaptability of MoE, it has been leveraged in several applications, namely speech recognition You_Speech_MoE ; You_Speech_MoE_2 , multi-task learning hazimeh_dselect_k_2021 , protein–protein interaction prediction (Qi2007, ), climate modeling (Lu2006, ), financial time-series prediction (Yumlu2003, ), and audio classification (Harb2004, ).
Subsequently, to upgrade the scalability of the standard MoE models, Shazeer et al. shazeer2017topk proposed a sparse MoE variant that selectively activates a limited subset of specialized experts for each input instance. This selective activation strategy significantly expands the model’s parameter capacity without a proportional increase in computational cost. Consequently, sparse MoE architectures have attracted widespread interest in a variety of fields such as natural language processing jiang2024mixtral ; Du_Glam_MoE ; fedus2022switch ; lepikhin_gshard_2021 , computer vision Riquelme2021scalingvision ; liang_m3vit_2022 , multimodal learning han2024fusemoe ; yun2024flex , and domain adaptation chi_representation_2022 ; nguyen2025cosine , where feed-forward layers in the Transformer architectures vaswani2017attention are substituted with sparse MoE layers.
Transformer models vaswani2017attention have recently become dominant in several applications, including language modeling devlin2019bert ; brown2020language ; raffel2020exploring ; touvron2023llama , computer vision dosovitskiy2020image ; carion2020end ; radford2021learning ; liu2021swin , and reinforcement learning chen2021decision ; kim2023preference ; hu2024learning . The core strength of the Transformers lies in their self-attention mechanism, which enables the models to dynamically focus on important parts of the input. Furthermore, self-attention also constructs context-sensitive representations for each token by weighting input elements according to their relevance, effectively capturing long-range and complex dependencies within the data. Mathematically, given an arbitrary input ${\bm{X}}\in\mathbb{R}^{N\times d}$ of $N$ feature vectors of dimension $d$, the self-attention mechanism first calculates its three key components, namely the query matrix ${\bm{Q}}\in\mathbb{R}^{N\times d_{k}}$, the key matrix ${\bm{K}}\in\mathbb{R}^{N\times d_{k}}$, and the value matrix ${\bm{V}}\in\mathbb{R}^{N\times d_{v}}$ as follows:
${\bm{Q}}={\bm{X}}\bm{W}_{q},\quad{\bm{K}}={\bm{X}}\bm{W}_{k},\quad{\bm{V}}={\bm{X}}\bm{W}_{v},$
where $\bm{W}_{q}\in\mathbb{R}^{d\times d_{k}},\bm{W}_{k}\in\mathbb{R}^{d\times d_{k}},\bm{W}_{v}\in\mathbb{R}^{d\times d_{v}}$ are trainable projection matrices. Next, attention scores, which measure the relevance of one element in an input sequence with respect to other elements, are computed as $\mathrm{Softmax}\Bigl{(}\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\Bigr{)}$, where the softmax function acts row-wise on the matrix ${\bm{Q}}{\bm{K}}^{\top}/\sqrt{d_{k}}\in\mathbb{R}^{N\times N}$. Then, the final output of the self-attention mechanism is given by
$\mathrm{Attn}({\bm{X}})=\mathrm{Softmax}\Bigl{(}\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\Bigr{)}{\bm{V}}=\mathrm{Softmax}\Bigl{(}\frac{{\bm{X}}\bm{W}_{q}\bm{W}_{k}^{\top}{\bm{X}}^{\top}}{\sqrt{d_{k}}}\Bigr{)}{\bm{X}}\bm{W}_{v}.$ (2)
From equations (1) and (2), it can be seen that both MoE and self-attention are written as a softmax weighted sum of functions of the input. Thus, there have been some attempts to explore the connection between these two popular models in the literature.
Recent work explores connections between self-attention and MoE frameworks by integrating MoE-inspired sparse routing into attention mechanisms and employing attention-based routing strategies within MoE models. SwitchHead (csordas2024switchhead, ) selectively routes queries to attention heads, significantly reducing computational and memory overhead, while Multi-Head MoE (wu2024multi, ) decouples experts from attention heads to enhance scalability. Mixture-of-Head (MoH) (jin2024moh, ) explicitly formalizes each head as an expert chosen by a learned router. Earlier approaches like Mixture of Attention Heads (MoA) (zhang2022moa, ) introduced per-token head mixtures as an alternative to uniform averaging, and Mixture of Softmax (MoS) (yang2018mos, ) aimed to overcome the expressiveness limitations of standard softmax attention distributions. MoEAtt (chen2023moeatt, ) further integrates attention-based routing with quadratic gating to provide greater flexibility in expert assignment. However, we notice that the aforementioned links between MoE and self-attention seem to be intuitive rather than rigorously derived.
@@
Table 1: Summary of expert estimation rates under MoE models equipped with the quadratic polynomial gate and the quadratic monomial gate. Below, $n$ stands for the sample size, while $\tau$ is some positive constant.
Gates FFN Experts Linear Experts
Quadratic Polynomial Gate $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(1/\log^{\tau}(n))$
Quadratic Monomial Gate $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(1/\log^{\tau}(n))$
Contributions. Therefore, in this paper, we aim to establish a stronger relation between these two models with theoretical guarantee. Furthermore, we also provide a comprehensive convergence analysis of the resulting MoE model from that relation to offer new insight into the design of self-attention. Our contributions are three-fold and can be summarized as follows:
1. Relation between self-attention mechanism and quadratic gating MoE. Building on the similarity in the formulations of self-attention and MoE in equations (1) and (2), we unify these two popular concepts by showing that each row of a self-attention matrix can be represented as a quadratic gating MoE (see Proposition 2.1). This relation motivates us to study the quadratic gating MoE in more detail.
@@
2. Theoretical analysis of the quadratic gating MoE. We explore the effects of two quadratic gating functions, namely the quadratic polynomial gating function and the quadratic monomial gating function, on the convergence rates of parameter estimation and expert estimation under the MoE models. For the convergence analysis of each gate, we provide a corresponding strong identifiability condition (see Definitions 3.2 and 3.6) to characterize the compatible structure of experts with that gate. These conditions indicate that approximating experts formulated as neural networks with non-linear activation functions require a much smaller amount of data than estimating linear experts (see Theorems 3.3 and 3.4).
@@
3. Practical implications. There are two important practical implications of the convergence analysis of quadratic gating MoE. Firstly, this analysis encourages the usage of non-linear experts over linear experts in quadratic gating MoE models (see Table 1). Given that insight, we propose a novel active-attention mechanism in equation (25) by applying a non-linear activation to the value matrix in the self-attention mechanism. Through extensive empirical evaluation, we show the favorable performance of the proposed active-attention over standard attention in various tasks, namely image classification, language modeling, and multivariate time series forecasting. Secondly, the theoretical results indicate that quadratic monomial gate yields faster expert and parameter estimation rates than quadratic polynomial gate (see Table 2). This property confirms the benefits of the widely used linear embeddings of the keys and queries without bias terms in the self-attention mechanism in practice.
Table 2: Summary of parameter estimation rates under the mixture of strongly identifiable experts equipped with the quadratic polynomial gate and the quadratic monomial gate. The function $\bar{r}$ below is defined in equation (14), while Voronoi cells $\mathcal{V}_{j}$ are defined in equation (16).
Gates $\exp(c^{*}_{j})$ ${\bm{A}}^{*}_{j}$ ${\bm{b}}^{*}_{j}$ ${\bm{\eta}}^{*}_{j}$
Quadratic Polynomial Gate $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{V}_{j}|)})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/2\bar{r}(|\mathcal{V}_{j}|)})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$
Quadratic Monomial Gate $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$ $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$
@@
Organization. The paper proceeds as follows. In Section 2, we establish a relation between the self-attention mechanism and MoE, and then introduce two quadratic gating functions of our interest. Next, we investigate the convergence behavior of parameter estimation and expert estimation under MoE models with these quadratic gates, following by two important practical implications of that analysis in Section 3. Then, we discuss related works to our paper in Section 4. After that, we conduct extensive experiments to empirically justify the theoretical results and favorable performance of the active-attention mechanism in Section 5. Then, we discuss related works to our paper in Section 4. Finally, we conclude the paper, highlight some limitations of our work, and list a few potential future directions in Section 6. Additional results and theoretical proofs are deferred to the Appendices.